Riverbend Inc. received a $327,500 dividend from stock it held in Hobble Corporation. Riverbend's taxable income is $2,200,000 before deducting the dividends received deduction (DRD), a $62,500 NOL carryover, and a $148,000 charitable contribution.

EXHIBIT 16-6 Stock Ownership and Dividends Received Deduction Percentage

Receiving Corporation’s Stock Ownership in Distributing Corporation’s Stock Dividends Received Deduction Percentage
Less than 20 percent    50%
At least 20 percent but less than 80 percent 65
80 percent or more* 100  

a. What is Riverbend’s deductible DRD assuming it owns 10 percent of Hobble Corporation?

b. Assuming the facts in part (a), what is Riverbend’s marginal tax rate on the dividend?

c. What is Riverbend’s DRD assuming it owns 76 percent of Hobble Corporation?

d. Assuming the facts in part (c), what is Riverbend’s marginal tax rate on the dividend?

e. What is Riverbend’s DRD assuming it owns 91 percent of Hobble Corporation (and is part of the same affiliated group)?

f. Assuming the facts in part (e), what is Riverbend’s marginal tax rate on the dividend?
Answer #1

Answers:

a. $163,750

Calculation

Riverbend owns 10% of Hobble which is less than 20 percent of Hobble, so its DRD percentage is 50%.

Its full DRD is $163,750 (0.50 × $327,500)

Riverbend's modified taxable income for the taxable income limitation is $2,052,000 ($2,200,000 minus $148,000 charitable contribution).

So, the taxable income limit =$1,026,000 ($2,052,000 × 50%).

As the full $163,750 DRD is less than the taxable income limit, Riverbend will deduct the entire $163,750 DRD.

b. 10.5%

Calculation

Corporate tax rate is 21%.

So, its effective tax rate on the dividend after taking the DRD into account is computed as follows:

($327,500 − $163,750) × 0.21] / $327,500 = 10.5%

c. $212,875

Riverbend owns 76% of Hobble, which is 20 percent or more but less than 80% of Hobble, so its DRD percentage is 65%.

Then, its full DRD = 0.65 × $327,500 = $212,875

Riverbend's modified taxable income for the taxable income limitation = $2,200,000 minus $148,000 (charitable contribution) = $2,052,000

so, the taxable income limit = $2,052,000 × 65% = $1,333,800.

As the full $212,875 DRD is less than the taxable income limit, Riverbend will deduct the entire $212,875 DRD.

d. 7.4%

Calculation

Corporate tax rate is 21%.

So, its effective tax rate on the dividend after taking the DRD into account is computed as follows:

($327,500 − $212,875) × 0.21] / $327,500 = 7.4%

e. $327,500

Calculation

Riverbend owns 91% of Hobble Corp which is 80% or more will be entitled to a 100% DRD.

f. 0%

Explanation

There will be no tax payable on dividend and So, the marginal tax rate of dividend is 0.

Riverbend does not pay any income tax on the dividend.